The percentage error in the measurement of mass of a body is 2 % and the percentage error in the measurement of the velocity is 0.75%. estimate in the percentage error in the momentum of the body

Grade:11

1 Answers

Vikas TU
4 years ago
Dear student 
% error in mass = 2% 
% error in velocity = 0.75 % 
momentum = m*v 
so error will add up , 
% error in momentum = 2+ 0.75 = 2.75 % 
Hope this helps 
Good luck